# UPGS: Unified Pose‑aware Gaussian Splatting for Dynamic Scene Deblurring

### v2

@@ Abstract
Reconstructing dynamic 3D scenes from monocular video has broad applications in AR/VR, robotics, and autonomous navigation, but often fails due to severe motion blur caused by camera and object motion. Existing methods commonly follow a two‑step pipeline, where camera poses are first estimated and then 3D Gaussians are optimized. Since blurring artifacts usually undermine pose estimation, pose errors could be accumulated to produce inferior reconstruction results. To address this issue, we introduce a unified optimization framework by incorporating camera poses as learnable parameters complementary to 3DGS attributes for end-to-end optimization. Specifically, we recast camera and object motion as per‑primitive SE(3) affine transformations on 3D Gaussians and formulate a unified optimization objective. For stable optimization, we introduce a three‑stage training schedule that optimizes camera poses and Gaussians alternatively. Particularly, 3D Gaussians are first trained with poses being fixed, and then poses are optimized with 3D Gaussians being untouched. Finally, all learnable parameters are optimized together. Extensive experiments on the Stereo Blur dataset and challenging real‑world sequences demonstrate that our method achieves significant gains in reconstruction quality and pose estimation accuracy over prior dynamic deblurring methods.
@@ 1 Introduction
Recent advances in explicit 3D Gaussian Splatting (3DGS) [8] and implicit Neural Radiance Fields (NeRF) [24] have pushed photorealistic reconstruction from static 3D scenes to fully dynamic 4D ones. These techniques promote emerging applications such as mobile AR/VR, autonomous driving, and robotics, in which data are almost acquired with a single, moving camera in unconstrained environments [22, 7, 50]. In these scenarios, severe motion-induced blur usually remains a dominant factor that limits reconstruction quality.
@@
Motion blur arises as the sensor integrates light over a finite exposure while the object or the camera moves, which consists of two components: camera-induced blur from ego-motion and object-induced blur from fast, often non-rigid motion [6, 31]. These motions produce blurring artifacts that hinder high-quality 4D reconstruction. The major reasons are twofold: (i) Low pose accuracy. Structure-from-motion pipelines such as COLMAP depend on sharp, repeatable keypoints to triangulate a sparse cloud and estimate camera poses. However, motion blur loses high-frequency details, frustrates feature matching, and produces inaccurate camera poses [30, 12]. (ii) High ill-posedness. As each blurred pixel integrates light from a trajectory rather than a fixed 3D location, which weakens the geometric constraint and increases the ill-posedness to reconstruct 4D scenes [10].
@@
Over the last years, numerous efforts [21, 15, 14, 39, 36, 13, 27] have been made to reconstruct sharp scenes from motion-blurred imagery, yet most methods still focus on static 3D scenes. Recently, dynamic deblurring has emerged [20, 10, 34, 2, 19] and is mainly built on NeRF. Since the continuous representation naturally favors smooth, low-frequency results, fast object motions are typically more difficult to be optimized, leading to unsatisfactory solutions. In addition, volumetric ray marching through a deep MLP requires long training time [24], making deployment for time-sensitive AR/VR use cases challenging. Recent advances in Gaussian splatting has provided a new approach, yet their applications in dynamic deblurring is still under-investigated.
@@
Intuitively, 4D reconstruction from a blurry video is a joint optimization problem in which camera pose estimation and sharp scene reconstruction are deeply coupled [17, 1]. On the one hand, sharp reconstruction improves the accuracy of pose estimation by providing higher-quality keypoints. On the other hand, accurate camera poses alleviate the ill-posedness to facilitate sharp images to be reconstructed. However, most prior approaches ignore the interaction between pose estimation and scene reconstruction. These methods commonly employ COLMAP poses to reconstruct the scene without further pose refinement. As a result, pose errors can be accumulated to degrade the final reconstruction results.
@@
  • We demonstrate on the Stereo Blur dataset and challenging real‑world sequences that UPGS yields sharper reconstructions and more accurate camera‑pose estimates, outperforming prior dynamic deblurring methods.
@@ 2 Related Work
In this section, we first review previous 2D deblurring methods, and then present advanced novel view synthesis methods. Finally, we focus on discussing recent advances regarding deblurring dynmaic reconstruction.
@@
The goal of deblurring is to recover sharp imagery from blur caused by camera shake or object motion. Classical methods cast this as an inverse‑filtering problem with hand‑crafted priors [4, 9, 29, 23], while modern deep‑learning approaches train end‑to‑end multi‑scale, recurrent, attention‑ and adversarial‑based networks [25, 35, 45, 11] on large datasets (e.g GoPro [25], DeblurGAN [11]), enabling real‑time single‑image deblurring. Video deblurring further leverages temporal cues by aligning or fusing features across frames via optical flow, deformable convolutions, or recurrent/3D networks [33, 38, 28, 48, 26]. Despite these advances in 2D image/video space, such models lack explicit 3D scene geometry modeling. Our work formulates deblurring as a 4D reconstruction problem within the Gaussian Splatting framework.
@@ 2.2 Dynamic Novel View Synthesis
Gaussian Splatting (GS) [8] has emerged as a popular choice for modeling 3D scenes due to its explicit structure, high rendering efficiency, and real-time performance. Recent works have extended GS into 4D to handle dynamic scene reconstruction from monocular input. Early attempts, such as Dynamic 3D Gaussians [18], independently tracked each Gaussian across timestamps to capture scene motion, while Deformable-3D-Gaussians [43] introduced a deformation field via an MLP to learn time-dependent transformations of canonical Gaussians. 4D Gaussian Splatting [41, 44, 37] advanced this idea by directly lifting Gaussians into 4D space, jointly encoding spatial and temporal variations to enable high-quality real-time dynamic rendering. More recent developments, including Disentangled 4D Gaussian Splatting [3], significantly improve efficiency by factorizing spatial and temporal components, while benchmarking efforts [16] have analyzed the performance and brittleness of various monocular 4D GS pipelines. These milestones have collectively established GS as a competitive framework for 4D reconstruction, capable of modeling complex motions while maintaining real-time rendering capabilities. However, existing 4D GS approaches still rely on sharp, high-quality inputs and accurate camera poses to achieve good reconstruction fidelity, which limits their robustness in real-world scenarios involving motion blur. To address this, our method aims to reconstruct 4D scenes directly from blurry monocular inputs, jointly handling deblurring and dynamic scene modeling.
@@ 2.3 Deblurring Dynamic Reconstruction
Recently, a variety of works have been proposed to recover high-quality 3D reconstructions from blurred inputs. NeRF-based approaches [21, 36, 14, 15] treat blur as part of the forward model by learning spatially varying kernels or simulating exposure via camera-trajectory subframes. 3DGS-based static deblurring folds the same exposure model into an explicit pipeline, with Debluring 3DGS [13] as a pioneering example. BAD-Gaussians [47] and DeblurGS [27] share a workflow that jointly estimates exposure-time camera motion and optimizes Gaussian scene parameters by rendering virtual sharp subframes along a continuous SE(3) trajectory, averaging them to synthesize the observed blur, and driving a photometric loss against the input images. These approaches deliver high-quality results on static scenes but do not handle dynamics. Turning to dynamic 4D deblurring, [34, 2, 20] generates latent rays within the exposure and performs time-aware volume rendering. However, training often requires significantly longer runtimes, and the implicit MLP’s smoothness and capacity bias toward low-frequency content causes fast non-rigid motions to be underfit. Our approach fills this gap with a 4D Gaussian-splatting formulation that reconstructs directly from blurry monocular video while explicitly decoupling camera-induced blur via subframe warps along with the object motion, improving efficiency and preserving high-frequency structure in dynamic areas.
@@ 3.1.1 Gaussian Splatting Primer
1:initial scene $\theta=\{\theta_{\text{sta}},\theta_{\text{dym}}\}$, initial poses $\phi$, stage bounds $E_{1},E_{2}$, total epochs $E_{\max}$
2:Definitions:
@@
[…]
8:for epoch $e=0$ to $E_{\max}-1$ do
9:  for $t$ in minibatch do
10:   if $e<E_{1}$ then
11:     $L\leftarrow L_{\mathrm{total}}(t;\theta,\bot(\phi))$
12:     $\theta\leftarrow\mathcal{U}_{\theta}(L)$
13:   else if $e<E_{2}$ then
14:     $L\leftarrow L_{\mathrm{total}}(t;\bot(\theta),\phi)$
15:     $\phi\leftarrow\mathcal{U}_{\phi}(L)$
16:   else
17:     $L\leftarrow L_{\mathrm{total}}(t;\theta,\phi)$
18:     $\theta\leftarrow\mathcal{U}_{\theta}(L)$
19:     $\phi\leftarrow\mathcal{U}_{\phi}(L)$
@@ 3.1.2 Gaussian Deformation for Object Motion
Following Deblur4DGS [42], object deformation is modeled by $\mathrm{SE}(3)$ transforms $\{A_{t},E_{t}\}$ via Shape‐of‐Motion [37], i.e.
@@ 3.2 Unified Pose-Aware Framework
Our unified pose-aware framework is illustrated in Fig. 1. To make camera pose trainable within the 3DGS framework, we formulate the camera motion as a affine transformation of the scene under a given fixed pose. In this way, camera motion can be modeled with object motion using unified representations, which are then applied to Gaussian primitives. Given a blurry frame $B_{t}$ at timestamp $t$, we estimate the camera pose using CLOMAP and employ it to initialize the affine transformation $P_{t}\in\mathrm{SE}(3)$. To capture the camera trajectory during exposure time, we subdivide the exposure time into $N$ intervals such that the camera poses for subframes can be denoted as:
@@
Figure 2: Visual Comparison on Stereo Blur Dataset. The orange boxes highlights regions with intense dynamic motion and the blue boxes indicate purely static areas.
@@
leaving color $c_{i}$ and opacity $\alpha_{i}$ unchanged. As a result, by rendering the warped 3D Gaussians $\{\bm{G}^{\prime}\}$ from the fixed reference pose $P_{t}^{(n)}$ and averaging over all $m$ subframes, we faithfully mimic the true motion blur— the integral of the scene’s geometry under the camera’s continuous motion—directly in the geometry domain.
@@ 3.3 Optimization Strategy
Since pose and geometry are tightly coupled, a fully joint optimization can be highly ill‑posed and produce trivial solutions. Minor errors in camera pose estimation may skew the scene fit, which in turn misguides subsequent pose updates, leading to error accumulation and unsatisfactory reconstruction results. To address this, we partition the training phase into three successive stages governed by the epoch index $e$ and two boundary epochs $E_{1}$ and $E_{2}$. The training strategy is summarized in Algorithm 1.
  • For $e<E_{1}$, we optimize only the scene Gaussians $\{\bm{G}_{\mathrm{sta}},\,\bm{G}_{\mathrm{dym}}\}$ with hold camera poses fixed at their SfM initialization
  • Then, for $E_{1}\leq e<E_{2}$, we freeze all Gaussian parameters and refine only the camera‑motion learnables $\{P_{t}^{(m)}\}$.
  • Finally, for $e\geq E_{2}$, we jointly optimize both geometry and pose in an end‑to‑end manner.
@@
$L_{\mathrm{dym}}\bigl{(}\bm{G_{\mathrm{sta}}},\,\bm{G_{\mathrm{dym}}};P_{t}^{(m)}\bigr{)}=\bigl{\lVert}I\bigl{(}\bm{G_{\mathrm{sta}}},\,\bm{G_{\mathrm{dym}}};P_{t}^{(m)}\bigr{)}-I(t)\bigr{\rVert}^{2},$ (11)
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Training time $\downarrow$
DyBluRF 23.82 0.690 0.471 49
Shape-of-Motion 27.49 0.922 0.192 3.7
4DGS 27.39 0.801 0.339 0.8
Deblur4DGS 29.02 0.745 0.165 6.9
Ours 30.14 0.911 0.107 4.2
Table 1: Quantitative results on the Stereo Blur dataset. Red and yellow denote the best and second-best per column, respectively.

### v3

@@ Abstract
Reconstructing dynamic 3D scenes from monocular video has broad applications in AR/VR, robotics, and autonomous navigation, but often fails due to severe motion blur caused by camera and object motion. Existing methods commonly follow a two‑step pipeline, where camera poses are first estimated and then 3D Gaussians are optimized. Since blurring artifacts usually undermine pose estimation, pose errors could be accumulated to produce inferior reconstruction results. To address this issue, we introduce a unified optimization framework by incorporating camera poses as learnable parameters complementary to 3DGS attributes for end-to-end optimization. Specifically, we recast camera and object motion as per‑primitive SE(3) affine transformations on 3D Gaussians and formulate a unified optimization objective. For stable optimization, we introduce a three‑stage training schedule that optimizes camera poses and Gaussians alternatively. Particularly, 3D Gaussians are first trained with poses being fixed, and then poses are optimized with 3D Gaussians being untouched. Finally, all learnable parameters are optimized together. Extensive experiments on the Stereo Blur and BARD-GS dataset with challenging real‑world sequences demonstrate that our method achieves significant gains in reconstruction quality and pose estimation accuracy over prior dynamic deblurring methods.
@@ 1 Introduction
Recent advances in explicit 3D Gaussian Splatting (3DGS) [kerbl20233d] and implicit Neural Radiance Fields (NeRF) [mildenhall2021nerf] have pushed photorealistic reconstruction from static 3D scenes to fully dynamic 4D ones. These techniques promote emerging applications such as mobile AR/VR, autonomous driving, and robotics, in which data are almost acquired with a single, moving camera in unconstrained environments [matsuki2024gaussian, jiang2024vr, zhou2024drivinggaussian]. In these scenarios, severe motion-induced blur usually remains a dominant factor that limits reconstruction quality.
@@
Motion blur arises as the sensor integrates light over a finite exposure while the object or the camera moves, which consists of two components: camera-induced blur from ego-motion and object-induced blur from fast, often non-rigid motion [ji2024motion, rozumnyi2022motion]. These motions produce blurring artifacts that hinder high-quality 4D reconstruction. The major reasons are twofold: (i) Low pose accuracy. Structure-from-motion pipelines such as COLMAP depend on sharp, repeatable keypoints to triangulate a sparse cloud and estimate camera poses. However, motion blur loses high-frequency details, frustrates feature matching, and produces inaccurate camera poses [peng2023pdrf, lee2024sharp]. (ii) High ill-posedness. As each blurred pixel integrates light from a trajectory rather than a fixed 3D location, which weakens the geometric constraint and increases the ill-posedness to reconstruct 4D scenes [kumar2025dynamode].
@@
Over the last years, numerous efforts [ma2022deblur, lee2023exblurf, lee2023dp, wang2024mp, wang2023bad, lee2024deblurring, oh2024deblurgs] have been made to reconstruct sharp scenes from motion-blurred imagery, yet most methods still focus on static 3D scenes. Recently, dynamic deblurring has emerged [luthra2024deblur, kumar2025dynamode, sun2024dyblurf, bui2025moblurf, luo2024dynamic] and is mainly built on NeRF. Since the continuous representation naturally favors smooth, low-frequency results, fast object motions are typically more difficult to be optimized, leading to unsatisfactory solutions. In addition, volumetric ray marching through a deep MLP requires long training time [mildenhall2021nerf], making deployment for time-sensitive AR/VR use cases challenging. Recent advances in Gaussian splatting has provided a new approach, yet their applications in dynamic deblurring is still under-investigated.
@@
Intuitively, 4D reconstruction from a blurry video is a joint optimization problem in which camera pose estimation and sharp scene reconstruction are deeply coupled [lin2021barf, bian2023nope]. On the one hand, sharp reconstruction improves the accuracy of pose estimation by providing higher-quality keypoints. On the other hand, accurate camera poses alleviate the ill-posedness to facilitate sharp images to be reconstructed. However, most prior approaches ignore the interaction between pose estimation and scene reconstruction. These methods commonly employ COLMAP poses to reconstruct the scene without further pose refinement. As a result, pose errors can be accumulated to degrade the final reconstruction results.
@@
  • We demonstrate that UPGS yields sharp reconstructions and more accurate camera‑pose estimates on challenging real‑world sequences from the Stereo Blur [zhou2019davanet] and BARD‑GS datasets [lu2025bard], indicating its effectiveness for dynamic deblurring tasks.
@@ 2 Related Work
In this section, we first review previous 2D deblurring methods, and then present advanced novel view synthesis methods. Finally, we focus on discussing recent advances regarding deblurring dynamic reconstruction.
@@
The goal of deblurring is to recover sharp imagery from blur caused by camera shake or object motion. Classical methods cast this as an inverse‑filtering problem with hand‑crafted priors [fergus2006removing, krizhevsky2012advances, pan2016blind, michaeli2014blind], while modern deep‑learning approaches train end‑to‑end multi‑scale, recurrent, attention‑ and adversarial‑based networks [nah2017deep, tao2018scale, zamir2021multi, kupyn2018deblurgan] on large datasets (e.g GoPro [nah2017deep], DeblurGAN [kupyn2018deblurgan]), enabling real‑time single‑image deblurring. Video deblurring further leverages temporal cues by aligning or fusing features across frames via optical flow, deformable convolutions, or recurrent/3D networks [su2017deep, wang2019edvr, pan2020cascaded, zhou2019spatio, nah2019recurrent]. Despite these advances in 2D image/video space, such models lack explicit 3D scene geometry modeling. Our work formulates deblurring as a 4D reconstruction problem within the Gaussian Splatting framework.
@@ 2.2 Dynamic Novel View Synthesis
Gaussian Splatting (GS) [kerbl20233d] has emerged as a popular choice for modeling 3D scenes due to its explicit structure, high rendering efficiency, and real-time performance. Recent works have extended GS into 4D to handle dynamic scene reconstruction from monocular input. Early attempts, such as Dynamic 3D Gaussians [luiten2024dynamic], independently tracked each Gaussian across timestamps to capture scene motion, while Deformable-3D-Gaussians [yang2024deformable] introduced a deformation field via an MLP to learn time-dependent transformations of canonical Gaussians. 4D Gaussian Splatting [wu20244d, yang2023real, wang2024shape] advanced this idea by directly lifting Gaussians into 4D space, jointly encoding spatial and temporal variations to enable high-quality real-time dynamic rendering. More recent developments, including Disentangled 4D Gaussian Splatting [feng2025disentangled], significantly improve efficiency by factorizing spatial and temporal components, while benchmarking efforts [liang2024monocular] have analyzed the performance and brittleness of various monocular 4D GS pipelines. These milestones have collectively established GS as a competitive framework for 4D reconstruction, capable of modeling complex motions while maintaining real-time rendering capabilities. However, existing 4D GS approaches still rely on sharp, high-quality inputs and accurate camera poses to achieve good reconstruction fidelity, which limits their robustness in real-world scenarios involving motion blur. To address this, our method aims to reconstruct 4D scenes directly from blurry monocular inputs, jointly handling deblurring and dynamic scene modeling.
@@ 2.3 Deblurring Dynamic Reconstruction
Recently, a variety of works have been proposed to recover high-quality 3D reconstructions from blurred inputs. NeRF-based approaches [ma2022deblur, wang2023bad, lee2023dp, lee2023exblurf] treat blur as part of the forward model by learning spatially varying kernels or simulating exposure via camera-trajectory subframes. 3DGS-based static deblurring folds the same exposure model into an explicit pipeline, with Debluring 3DGS [lee2024deblurring] as a pioneering example. BAD-Gaussians [zhao2024bad] and DeblurGS [oh2024deblurgs] share a workflow that jointly estimates exposure-time camera motion and optimizes Gaussian scene parameters by rendering virtual sharp subframes along a continuous SE(3) trajectory, averaging them to synthesize the observed blur, and driving a photometric loss against the input images. These approaches deliver high-quality results on static scenes but do not handle dynamics. Turning to dynamic 4D deblurring, [sun2024dyblurf, bui2025moblurf, luthra2024deblur] generates latent rays within the exposure and performs time-aware volume rendering. However, training often requires significantly longer runtimes, and the implicit MLP’s smoothness and capacity bias toward low-frequency content causes fast non-rigid motions to be underfit. Our approach fills this gap with a 4D Gaussian-splatting formulation that reconstructs directly from blurry monocular video while explicitly decoupling camera-induced blur via subframe warps along with the object motion, improving efficiency and preserving high-frequency structure in dynamic areas.
@@ 3.1.1 Gaussian Splatting Primer
1:scene $\theta$, poses $\phi$, epochs $E_{\max}$, stop-gradient operator $\bot(\cdot)$
@@
3:for $e=0$ to $E_{\max}-1$ do
4:  for $t$ in minibatch do
5:   if $\text{stage}=1$ then $\triangleright$ scene warm-up
6:     $\theta\leftarrow\mathcal{U}_{\theta}\!\big(L_{\text{total}}(t;\theta,\bot(\phi))\big)$
7:     if $\Delta_{\text{psnr}}<\epsilon_{1}\wedge\text{anchor}>\tau_{a}$ then
8:      $\text{stage}\leftarrow 2$
9:     end if
10:   else if $\text{stage}=2$ then $\triangleright$ pose refinement
11:     $\phi\leftarrow\mathcal{U}_{\phi}\!\big(L_{\text{total}}(t;\bot(\theta),\phi)\big)$
12:     if $\Delta_{\text{pose}}<\epsilon_{2}\wedge\mathcal{P}[\mathcal{L}]>\tau_{\text{gain}}$ then
13:      $\text{stage}\leftarrow 3$
14:     end if
@@
16:     $\theta\leftarrow\mathcal{U}_{\theta}\!\big(L_{\text{total}}(t;\theta,\phi)\big)$;   $\phi\leftarrow\mathcal{U}_{\phi}\!\big(L_{\text{total}}(t;\theta,\phi)\big)$
17:   end if
18:  end for
19:end for
@@ 3.1.2 Gaussian Deformation for Object Motion
Following Deblur4DGS [wu2024deblur4dgs], object deformation is modeled by $\mathrm{SE}(3)$ transforms $\{A_{t},E_{t}\}$ via Shape‐of‐Motion [wang2024shape], i.e.
@@ 3.2 Unified Pose-Aware Framework
Our unified pose-aware framework is illustrated in Fig. 1. To make camera pose trainable within the 3DGS framework, we formulate the camera motion as an affine transformation of the scene under a given fixed pose. In this way, camera motion can be modeled with object motion using unified representations, which are then applied to Gaussian primitives. Given a blurry frame $B_{t}$ at timestamp $t$, we estimate the camera pose using CLOMAP and employ it to initialize the affine transformation $P_{t}\in\mathrm{SE}(3)$. To capture the camera trajectory during exposure time, we subdivide the exposure time into $N$ intervals such that the camera poses for subframes can be denoted as:
@@
Figure 2: Visual Comparison on Stereo Blur and BARD-GS Dataset. The orange boxes highlights regions with intense dynamic motion and the blue boxes indicate purely static areas.
@@
leaving color $c_{i}$ and opacity $\alpha_{i}$ unchanged. As a result, by rendering the warped 3D Gaussians $\{\bm{G}^{\prime}\}$ from the fixed reference pose $P_{t}^{(n)}$ and averaging over all intermediate subframes, we faithfully mimic the true motion blur— the integral of the scene’s geometry under the camera’s continuous motion—directly in the geometry domain.
@@ 3.3 Optimization Strategy
Since pose and geometry are tightly coupled, a fully joint optimization can be highly ill‑posed and produce trivial solutions. Minor errors in camera pose estimation may skew the scene fit, which in turn misguides subsequent pose updates, leading to error accumulation and unsatisfactory reconstruction results. To address this, we partition the training phase into three successive stages: Stage 1 optimizes only the scene parameters, warm up the model with the basic scene geometry. Stage 2 freezes the scene and updates only the camera poses, enabling the system to digest the camera motion and understand the source of blur. Stage 3 then performs full joint optimization of both scene and pose. The overall training strategy is summarized in Algorithm 1.
The transition between stages is determined automatically by two adaptive gates. Gate 1 monitors whether the scene representation has saturated. It checks:
  • $\Delta_{\mathrm{psnr}}<\epsilon_{1}$: whether the increase in validation PSNR has plateaued.
  • $\mathrm{anchor}>\tau_{a}$: the extent to which the static Gaussians begin to shift as the model attempts to explain residual errors by implicitly mimicking camera motion.
$\text{anchor}=\sum_{g\in G_{\text{static}}}\left(\lVert\Delta\mu_{g}\rVert^{2}+\lVert\Delta R_{g}\rVert^{2}\right)$ (11)
Once these criteria are met, the scene has extracted the majority of learnable appearance structure, and pose optimization is activated.
Gate 2 evaluates whether pose refinement is exhausted. It detects:
  • $\Delta_{\mathrm{pose}}<\epsilon_{2}$: whether the decrease in pose error has plateaued.
  • $\mathcal{P}[\mathcal{L}](\theta,\phi)>\tau_{\text{gain}}$: whether enabling scene updates would yield a meaningful improvement in the overall objective. To assess this, we simulate a virtual optimizer step on the scene parameters as if scene learning were enabled. We then evaluate the corresponding decrease in loss by computing a lightweight look-ahead gain proxy from the scene gradients. A marked improvement is identified when incorporating scene parameters results in a clear reduction in total loss.
$\mathcal{P}[\mathcal{L}](\theta,\phi)=L_{\mathrm{pose}}(\phi)-L_{\mathrm{joint}}(\theta,\phi)$ (12)
When both conditions hold, the optimizer proceeds to Stage 3 for joint refinement of geometry and camera motion. The detailed values for the gating hyperparameters are provided in the supplementary material.
Table 1: Quantitative results on the Stereo and BARD datasets. Red and yellow denote the best and second-best per column, respectively.
Stereo Blur Dataset BARD Dataset
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Training time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Training time $\downarrow$
DyBluRF 23.82 0.690 0.471 49 16.25 0.523 0.971 87
Shape-of-Motion 27.49 0.922 0.192 3.7 17.92 0.718 0.434 8
4DGS 26.33 0.733 0.345 0.8 17.28 0.609 0.877 1.8
Deblur4DGS 29.02 0.745 0.165 6.9 18.80 0.720 0.372 13
Ours 30.14 0.911 0.107 4.2 19.50 0.731 0.550 10
@@
$L_{\mathrm{dym}}\bigl(\bm{G_{\mathrm{sta}}},\,\bm{G_{\mathrm{dym}}};P_{t}^{(m)}\bigr)$ (13)
$=$ $\bigl\lVert I\bigl(\bm{G_{\mathrm{sta}}},\,\bm{G_{\mathrm{dym}}};P_{t}^{(m)}\bigr)-I(t)\bigr\rVert^{2},$
@@
3.4 Discussion
We position UPGS against existing joint scene-pose deblurring pipelines by highlighting how it departs from prior neural-based approaches. BARF [lin2021barf] resolves pose–scene entanglement by gradually unmasking high-frequency positional encodings inside NeRF. However, it applies only to sharp, static input. We perform joint optimisation and reconstruction from dynamic, motion-blurred video. Unlike Deblur4DGS [wu2024deblur4dgs] which first optimises a frozen static background and then deblurs the dynamic foreground with extra regularisation, UPGS trains static and dynamic Gaussians together from the outset, avoiding additional loss terms. The recent BARD-GS [lu2025bard] learns one global pose per frame in camera space and interpolates intermediate viewpoints, while we represents both scene and pose in a unified geometry space and directly captures the entire camera trajectory.